# The Riemann Hypothesis is false

### v31

@@
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{2}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
@@
for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{4}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
@@
Let $\varepsilon=\varepsilon_{T}=T^{-1/4}$. We split the sum on the right-hand side of (9) for $m\leq\frac{n}{1+\varepsilon},\frac{n}{1+\varepsilon}\leq m\leq n-1,n+1\leq m\leq n(1+\varepsilon)$ and $m>n(1+\varepsilon)$. For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$ so it follows from (5) by routine partial summation (Riemann-Stieltjies integration) that
@@
It now remains to treat the ranges $\frac{m}{1+\varepsilon}\leq n\leq m-1$ and $m+1\leq n\leq m(1+\varepsilon)$, where $n\geq\frac{T}{\log^{2}T}$. Note that
$\sum_{\frac{m}{1+\varepsilon}<n<m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}<n<m-1}\frac{1}{m|\log(n/m)|}\ll 1$ (12)
@@
$\sum_{m+1<n<m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1<n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll 1.$ (13)
Notice that (13) is obtained by applying the fact that
$\sum_{m+1<n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\frac{1}{m}\int_{m+1}^{m(1+\varepsilon)}\frac{\mathrm{d}x}{\log(x/m)}\ll\frac{x/m}{m\log(x/m)}\Bigg{|}_{m+1}^{m(1+\varepsilon)}\ll 1.$ (14)
A similar argument yields (12). Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (13), (12), (11) and (10) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields

### v32

@@
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
for $\sigma\in\mathbb{R},|\sigma|\leq 2$. Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
@@
for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{-1/2}\log^{8}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
@@
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$. For $m\geq T(1+\varepsilon)$ and $m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$. Thus it follows from (5) by routine partial summation (Riemann-Stieltjes integration) that for such $m$ we have
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (10)
For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it follows again from (5) by partial summation that
@@
We now treat the ranges $\frac{m}{1+\varepsilon}<n\leq m-1$ and $m+1\leq n<m(1+\varepsilon)$. Note that for $m\ll T$, we have
$\sum_{\frac{m}{1+\varepsilon}<n\leq m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}<n\leq m-1}\frac{1}{m|\log(n/m)|}\ll\log T$ (13)
@@
$\sum_{m+1\leq n<m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1\leq n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\log T.$ (14)
Notice that (14) is obtained by writing $n=m+r$ for $11\leq r\leq\varepsilon m-$ 1 and applying the fact that $\log(1+\delta)\sim\delta$ uniformly for $1/m\leq\delta\leq\varepsilon$, hence
$\sum_{m+1\leq n<m(1+\varepsilon)}\frac{1}{m\log(n/m)}\ll\sum_{1\leq r\leq\varepsilon m-1}\frac{1}{r}\ll\log T.$ (15)
A similar argument yields (13). Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (14), (13), (12), (11) and (10) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields